OpenStudy (anonymous):

Solve for x: 3x - 5 = 2x + 6. 1 -1 11 -11

OpenStudy (anonymous):

how to do it

OpenStudy (anonymous):

Subtract 2x from each side

OpenStudy (anonymous):

Then add 5 to each side.

OpenStudy (anonymous):

3x-5=2x+6 3x=2x+11 x=11

OpenStudy (texaschic101):

we are going to get our x terms on one side and everything else on the other side. And remember, what you do on one side of the equal sign, you have to do on the other side. 3x - 5 = 2x + 6 -- subtract 2x from both sides 3x - 2x - 5 = 6 -- add 5 to both sides 3x - 2x = 6 + 5 -- combine like terms x = 11 check.. 3x - 5 = 2x + 6 3(11) - 5 = 2(11) + 6 33 - 5 = 22 + 6 28 = 28 (correct) x = 11 any questions ?

OpenStudy (anonymous):

no thanks guys

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

ur welcome.